OpenStudy (lonnie2018):

put x – 2y = 6 in slope intercept form

OpenStudy (unklerhaukus):

slope-intercept form is y = mx + b where m is the slope, and b is the y-intercept.

OpenStudy (unklerhaukus):

so solve the equation for y,

OpenStudy (shadowlegendx):

I already know you got this ;P

OpenStudy (anonymous):

\[x-2y=6\\ x-6=2y\\ \implies\ 2y=x-6\\ \implies\ y=\frac{ x-6 }{ 2 }\\ \implies\ y= \frac{ 1 }{ 2}x-\frac{ 6 }{ 2 }\\ \implies\ y=\frac{ 1 }{ 2}x-3\\ which \ is\ desired\ slope \ intercept\ form\ of\ given \ linear\ equation.\]